OpenStudy (errinss):

A drinking glass is shaped like a cylinder with a height of 12 cm and a radius of 4 cm. Maeva adds 25 spherical pieces of ice to the glass. The pieces of ice each have a diameter of 3 cm. How many cubic centimeters of water can Maeva add to fill the glass to the rim? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (sweetburger):

So first you want to find the volume of the drinking glass. Then you want to add up the area of the 25 spheres of ice. Subtract the volume of the combine 25 spheres from the volume of the glass. The resulting value is the amount of liquid you can add to to the glass.

OpenStudy (errinss):

uhm..... im lost to many words, can you simplify that?

OpenStudy (errinss):

@sweetburger

OpenStudy (sweetburger):

First fine the Volume of the drinking glass.

OpenStudy (sweetburger):

find*

OpenStudy (emmily_cool):

The volume of the cylinder is L* \pi r ^{2} Which is 12 * 3.14 * 4 ^{2} =602.88 cubic cm Each piece of ice's volume is defined as 4/3* \pi *r ^{3} cubic cm (volume of a sphere). With r=3/2 (since 3 cm is the diameter) =4/3 * 3.14 * (3/2) ^{3} =14.13 cubic cm 25 of those cumulate a total volume of 25 * 14.13 = 353.25 cm ^{3} Which means that the free space for water is 602.88-353.25=249.63 cubic cm

OpenStudy (sweetburger):

yahoo answers op lol

OpenStudy (errinss):

so i guess the answer is 249.63, right?